OpenStudy (anonymous):

A circular oil spill on the surface of the ocean spreads outward. Find the approximate rate of change in area of the oil slick with respect to its radius when the radius is 100m.

OpenStudy (anonymous):

The answer is 200pi in m^2/m units btw.

OpenStudy (dumbcow):

A = pi*r^2 dA/dr = 2*pi*r

OpenStudy (anonymous):

What does dA and dR mean?

OpenStudy (dumbcow):

derivative of function A with respect to variable r

OpenStudy (dumbcow):

no different than finding derivative of pi*x^2 which is 2*pi*x